OpenStudy (anonymous):

A barrel like a cylinder is laid on its side and rolled up a ramp that is 162ft long. The barrel has a circular base that is 1.1ft in diamter. How many times does the barrel turn in being rolled up the ramp? Use the valuse 3.14 for "pie". Round your answer to the neares tenth. Do not round any intermediate steps.

OpenStudy (anonymous):

Look for the circumference of the barrel. C = 2pi*r C = 2(3.14)(1.1/2) Then, divide the length of the ramp with the circumference of the barrel. The answer is the number of turns.

OpenStudy (anonymous):

\[162/(\pi 1.1) = 56.7 \]

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

Your welcome